OpenStudy (anonymous):

Please help & Explain! fg-9h=1oj (Solve for g)

jimthompson5910 (jim_thompson5910):

your first step is to move that "-9h" term over you do this by adding 9h to both sides (since addition undoes subtraction) to get \[\large fg-9h=1oj\] \[\large fg-9h+9h=1oj+9h\] \[\large fg+0h=1oj+9h\] \[\large fg+0=1oj+9h\] \[\large fg=1oj+9h\]

jimthompson5910 (jim_thompson5910):

what's your next step?

OpenStudy (anonymous):

wait im confused in step 3 where di you get "0" from?

jimthompson5910 (jim_thompson5910):

-9h + 9h = 0h since -9 + 9 = 0

jimthompson5910 (jim_thompson5910):

and because they add to 0, the "9h" terms go away on the left side (which is what we want) does that make sense

OpenStudy (anonymous):

ok i get that now. i dont no what to do next

jimthompson5910 (jim_thompson5910):

fg means "f times g"

jimthompson5910 (jim_thompson5910):

what is the opposite of multiplication?

OpenStudy (anonymous):

division

jimthompson5910 (jim_thompson5910):

so that means you divide both sides by f to isolate g \[\large fg=1oj+9h\] \[\large \frac{fg}{f}=\frac{1oj+9h}{f}\] \[\large \frac{1g}{1}=\frac{1oj+9h}{f}\] \[\large 1g = \frac{1oj+9h}{f}\] \[\large g = \frac{1oj+9h}{f}\]

jimthompson5910 (jim_thompson5910):

hopefully all that makes sense